OpenStudy (anonymous):

3x-9y =81 2x-6y=-10 solve then classify the system

OpenStudy (dumbcow):

did you try elimination method or addition method

OpenStudy (anonymous):

ive been working with the elimination method i wonder if the substitution method might work too but i havent gotten the right answer either way

OpenStudy (dumbcow):

either way will work for this problem i think elimination is easier lets cancel out the x's LCM of 2 and 3 is 6 multiply 1st row by -2 multiply 2nd row by 3 -6x +18y = -162 6x -18y = -30 ______________________ 0 + 0 = -192 0 = -192 this is False No solution

OpenStudy (dumbcow):

system is inconsistent

OpenStudy (anonymous):

if this occurs, then it's Undefined...

OpenStudy (anonymous):

that means it would be dependent as well right?

OpenStudy (anonymous):

x=dependent y=independent and since u don't know either one. "Undefined" then it can't be dependent or independent. unless i'm not seeing something here.

OpenStudy (dumbcow):

what do you mean undefined?? no not dependent, thats only if there are infinite solutions no solution means these 2 lines are parallel, same slope

OpenStudy (anonymous):

Oh. True sorry i wasn't looking at it like this. Just haven't done this for a long time...

OpenStudy (anonymous):

ok so it would be independent and inconsistent since both are parallel

OpenStudy (dumbcow):

correct

OpenStudy (anonymous):

ok now the last part of the question to this one is to show it in graph form what would this problem look like graphed

OpenStudy (dumbcow):

where do you think you were messing up?

OpenStudy (dumbcow):

graph both lines either by putting them in y=mx+b form or make a table of x,y values and connect the points

OpenStudy (anonymous):

whats bad is im trying to learn algebra in a 9 week period and ive never done it in my life back in high school we had the whole yr to learn this

OpenStudy (dumbcow):

oh online class ? yeah that can be hard

OpenStudy (anonymous):

it is very frustrating

OpenStudy (anonymous):

so y=(81-3x)/-9 and y=(-10-2x)/-6

OpenStudy (anonymous):

I've taken online math before. But it's always gave me an option to view and example of the problem how to solve. It's harder for u...

OpenStudy (anonymous):

im posting problems i have trouble with and am hoping to see them worked out with the answer so i can see how they are done but many just give me the answer which is not helping

OpenStudy (anonymous):

the graphing problems are hardest becaise noone really can show me what they look like

OpenStudy (anonymous):

have you herd of Graphing Calculators? Texas Instruments...

OpenStudy (dumbcow):

@beehive yeah those examples can be really helpful but they dont answer everyones questions and you dont know why you do the steps it shows you to do

OpenStudy (dumbcow):

graphcalc.com free graphing software

OpenStudy (anonymous):

ill chk it out im gonna post a few more and head to bed my head istarting to hurt lol